OpenStudy (anonymous):

Write an equation for the line that contains the point (-2, 9) and is parallel to the line y = 2x + 6. A. y = -(1/2)x + 8 B. y = 2x – 7 C. y = 2x + 11 D. y = 2x + 13

OpenStudy (anonymous):

Since you have a point and a slope, you can use point slope form. Do you know how to do this?

OpenStudy (anonymous):

y - y1 = m(x - x1) where m = slope

OpenStudy (anonymous):

Hello????

OpenStudy (anonymous):

yeah hold up im trying to do the equation on paper

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

would the answer be c

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.